OpenStudy (anonymous):

estimate cuberoot 30 using local linear approximation

OpenStudy (anonymous):

Okay, so \[ 30 = x^2 \]We want to find the roots of: \[ f(x) = x^2-30 \]

OpenStudy (anonymous):

no its cube root not square root

OpenStudy (anonymous):

Okay, well \(3^3 = 27\) and \(4^3 = 64\) so clearly \(27\) is closest cube.\[ f(x) = x^{1/3}\implies df =\frac 13 x^{-2/3}dx \]We use use the fact that when \(dx=\Delta x = x-a\): \[ \Delta f \approx df\implies f(x)-f(a)\approx \frac{1}{3}(a)^{-2/3}(x-a) \]

OpenStudy (anonymous):

In this case \(x=30\) and \(a=27\).

OpenStudy (anonymous):

\[ f(30)-f(27)\approx \frac{1}{3}(27)^{-2/3}(30-27) \]